# Excitable Nonlinear Opinion Dynamics (E-NOD) for Agile Decision-Making

### v2

@@ Abstract
We present Excitable Nonlinear Opinion Dynamics (E-NOD), which describe opinion-forming and decision-making behavior with superior “agility” in responding and adapting to fast and unpredictable changes in context, environment, or information received about available options. E-NOD derives through the introduction of a single extra term to the previously presented Nonlinear Opinion Dynamics (NOD), which have been shown to provide fast and flexible multi-agent behavior. The extra term is inspired by the fast-positive, slow-negative mixed-feedback structure of excitable systems. The agile behaviors brought about by the new excitable nature of decision-making driven by E-NOD are analyzed in a general setting and illustrated in an application to robot navigation around human movers.
@@ I Introduction
The nonlinear opinion dynamics (NOD) presented in [1, 2] model the time evolution of opinions of a group of agents engaged in a collective decision-making process over a set of options. The derivation of NOD was tailored to model and study the principles of adaptive decision-making in biological [3, 4, 5, 6] collectives and to use these principles to design adaptive decision-making in built collectives [1, 7, 8]. From a systems perspective, NOD exhibits a mixed-feedback [9, 10] structure: opinion formation arises from the balance of a negative feedback loop that regulates the agents’ opinions to a neutral state and many positive feedback loops (both at the single-agent and network levels) that destabilize the neutral state and trigger nonlinear opinion formation.
Decision-making driven by NOD is fast and flexible [2]. It is fast because it can exhibit fast divergence from indecision even in the absence of informative inputs about the options. It is flexible because in the presence of informative inputs the sensitivity of opinion formation to the information the inputs carry is tunable and adaptive to context. Both speed and flexibility of decision-making driven by NOD are determined by the existence of a tunable threshold for opinion formation. At this threshold, negative and positive feedback are perfectly balanced and the dynamics become singular.
@@
Figure 1: (a) Trajectory of a robot controlled with NOD (E-NOD) is shown with a blue (pink) line as it navigates towards a goal (star) in the presence of an oncoming human mover (black). The NOD robot experiences a collision; the E-NOD robot does not. (b) Opinion $z$ of the robot over time $t$. Circles denote matching time points along trajectories and opinions.
Figures 1, 4, and 5 animated in video here.
@@
Here, we introduce a single extra term in the NOD equations that enables the following “agile” decision-making behaviors: a) autonomous sequential decision-making, not requiring any ad-hoc reset of the model state once a decision is made; b) fast “changes of mind” and adaptive responses, when the context or information about the options change rapidly and unexpectedly; c) on-demand, or event-based, opinion formation in the sense that large opinions are formed sparsely in time in the form of “decision events” and only when context requires it; d) low energy and low resource use for greedy decision-making, resulting from the dynamics being event-based. These new agile behaviors are illustrated in Fig. 1 in the context of a robot navigating around a human mover as studied in [11].
The term we introduce in NOD is a slow regulation term inspired by the dynamics of excitable (spiking) signal processing systems, widespread in nature [12, 13, 14] and foundational to spiky control systems [15] and neuromorphic engineering [16]. The result is event-based decision-making that is a form of multi-dimensional excitability in the opinion formation space, i.e., spiking can occur in any of the multiple directions corresponding to the multiple options. This multi-dimensional excitability generalizes the one-dimensional excitability of classical spiking systems. We call NOD with the extra regulation term Excitable Nonlinear Opinion Dynamics (E-NOD).
@@
Our contributions are fivefold. First, in Section II, we present a new analysis of the singularity in NOD for a single decision-making agent and two options. We prove how a feedback gain, $K_{u}$, tunes opinion formation. We show that when $K_{u}$ gets too large, an opinion can become so robust that it will not change quickly enough if a new input arrives in favor of the alternative option (blue in Fig. 1). Second, in Section III, we present E-NOD for a single agent and two options. We show existence of the limit cycles associated with excitable behavior. We show geometrically the agility in behavior (pink in Fig. 1). Fourth, we generalize E-NOD to a network of multiple agents (Section IV). Fifth, we apply E-NOD to the social robot navigation problem (Section V).
II Fast and flexible decision-making
@@
We recall in Section II-A the NOD [1, 2] for a single agent evolving continuously over time its real-valued opinion about two mutually exclusive options in the possible presence of input. In Section II-B we analyze stability of the neutral solution and prove conditions on feedback gain $K_{u}$ that determine the type of singularity (type of pitchfork bifurcation) in the dynamics. We show that, by shaping bifurcation branches, $K_{u}$ tunes opinion formation and enables agile decision-making. In Section II-C, we show limits on tunability of NOD, which motivates introducing E-NOD in Section III.
II-A NOD for Single Decision-Maker
@@
We let an agent represent a single decision-maker. Let $z(t)\in\mathbb{R}$ define the agent’s opinion at time $t$ about two mutually exclusive options. The more positive (negative) is $z$, the more the agent favors (disfavors) option 1 and disfavors (favors) option 2. When $z\!=\!0$ the agent is neutral about the two options, i.e., in a state of indecision. Let $u(t)\!\geq\!0$ define the attention of the agent at time $t$ to its observations; $u$ is implemented as a gain in the dynamics. Let $b(t)\in\mathbb{R}$ define an input signal at time $t$ that represents external stimulus and/or internal bias. When $b(t)\!>\!0$ ($b(t)\!<\!0$), it provides information (evidence) in favor of option 1 (option 2).
@@
Decision-making variables $z$ and $u$ evolve continuously over time $t$ according to the following NOD, adapted from [1, 2]:
$\tau_{z}\dot{z}$ $=-d\,z+\tanh\big{(}u\left(\alpha z\right)+b\big{)}$ (1a)
@@
where ${\dot{z}}:=\!dz/dt$. $\tau_{z}\!>\!0$ is a time constant, and damping coefficient $d\!>\!0$ weights the negative feedback on $z$ that regulates the neutral state $z\!=\!0$. The second term in (1a) provides a nonlinear positive feedback on $z$ with weight given by the product of $u$ and amplification coefficient $\alpha\!>\!0$, plus the effects of $b$. The saturation nonlinearity given by the $\tanh$ function enables fast-and-flexible decision-making through opinion-forming bifurcations [1, 2]. The positive feedback gain is state-dependent according to (1b) and grows with $z^{2}$. Hence, small deviations from the neutral state in response to small inputs leave attention low and do not trigger large, nonlinear opinion formation. Large enough deviations from the neutral state in response to large enough inputs cause a sharp increase in attention and trigger large, nonlinear opinion formation. The resulting implicit threshold distinguishing small and large inputs is tuned by parameters $u_{0}$ and $K_{u}$.
II-B Analysis of NOD
We study the dynamics and stability of solutions of system (1a)-(1b) using bifurcation theory. A local bifurcation refers to a change in number and/or stability of the equilibrium solutions to a nonlinear dynamical system as a (bifurcation) parameter is changed. The state and parameter values at which this change occurs is called a bifurcation point. At a bifurcation point, one or more of the eigenvalues of the Jacobian of the model must have zero real part [17, 18], i.e., a bifurcation point is a singularity of the model vector field.
We are mainly interested in the pitchfork bifurcations. There are two generic types of pitchforks. A supercritical pitchfork bifurcation describes how one stable solution becomes unstable and two stable solutions emerge as the bifurcation parameter increases. A subcritical pitchfork bifurcation describes how two unstable solutions disappear and one stable solution becomes unstable as the bifurcation parameter increases.
@@ II-B Analysis of NOD
$\tau_{z}\dot{z}=-d\,z+\tanh\Big{(}(u_{0}+K_{u}z^{2})\cdot\left(\alpha z\right)+b\Big{)}.$ (2)
@@
$\dot{z}=\frac{1}{\tau_{z}}\Bigg{(}\left(\alpha\,u_{0}-d\right)z+\alpha\left(K_{u}-\frac{\alpha^{2}u_{0}^{3}}{3}\right)z^{3}\\ +\alpha^{3}u_{0}^{2}\left(\frac{2\alpha^{2}u_{0}^{3}}{15}-K_{u}\right)z^{5}\Bigg{)}+\mathcal{O}(z^{7}).$ (3)
A singularity exists at $(u_{0},z)\!=\!(u_{0}^{*},0)$, with $u_{0}^{*}\!=\!\frac{d}{\alpha}$. The solution $z\!=\!0$ is stable (unstable) when $u_{0}\!<\!u_{0}^{*}$ ($u_{0}\!>\!u_{0}^{*}$).
@@
When $z\!=\!0$ and $b\!=\!0$ the righthand side of (2) is zero so $z\!=\!0$ is always an equilibrium. We expand (2) with $b\!=\!0$ about $z\!=\!0$. The Taylor expansion of the hyperbolic tangent is $\tanh(w)\!=\!w-w^{3}/3+2w^{5}/15+\mathcal{O}(w^{7})$. Substituting this into (2) yields (3). The Jacobian $J\!=\!\frac{d\dot{z}}{dz}$ of (3) evaluated at $z\!=\!0$ is $J(0)\!=\!(\alpha\,u_{0}-d)/\tau_{z}$, which is singular when $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{\alpha}$. When $u\!<\!u_{0}^{*}$, $J(0)\!<\!0$ so $z\!=\!0$ is exponentially stable and when $u\!>\!u_{0}^{*}$, $J(0)\!>\!0$ so $z\!=\!0$ is unstable. ∎
@@
($K_{u}$ determines the singularity type): Consider (2). Let $b\!=\!0$ and $u_{0}^{*}\!=\!\frac{d}{\alpha}$. The singularity of dynamics (2) at $(u,z)\!=\!(u_{0}^{*},0)$ as proved in Lemma 1 corresponds to a supercritical pitchfork bifurcation for $K_{u}\!<\!\frac{d^{3}}{3\alpha}$, a quintic pitchfork bifurcation for $K_{u}\!=\!\frac{d^{3}}{3\alpha}$, and a subcritical pitchfork bifurcation for $K_{u}\!>\!\frac{d^{3}}{3\alpha}$.
@@
We define cubic coefficient $p$ to be the coefficient of $\alpha z^{3}/\tau_{z}$ in (3) at $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{\alpha}$. Then $p\!=\!(K_{u}\!-\!\frac{d^{3}}{3\alpha})$ and $p\!<\!0$ ($p\!>\!0$) for all $K_{u}\!<\!\frac{d^{3}}{3\alpha}$ ($K_{u}\!>\!\frac{d^{3}}{3\alpha}$). When $p\!<\!0$ ($p\!>\!0$), the term stabilizes (destabilizes) the system and produces a supercritical (subcritical) pitchfork bifurcation. We define quintic coefficient $q$ to be the coefficient of $(\alpha^{3}u_{0}^{2})z^{5}/\tau_{z}$ in (3) at $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{\alpha}$. Then $q\!=\!(\frac{2d^{3}}{15\alpha}\!-\!K_{u})$. Coefficients $p$ and $q$ as functions of $K_{u}$ are plotted in Fig. 2b. The results follow by invoking the recognition problem for the quintic pitchfork [17, Prop. VI.2.14] and the form of its $Z_{2}$-symmetric universal unfolding [17, Prop. VI.3.4 and Fig. VI.3.3]. ∎
Figure 2: The effect of $K_{u}$ on the bifurcation diagram of (2) and the cubic and quintic terms of (3). (a): Bifurcation diagrams of NOD (2) with $K_{u}$ values corresponding to the vertical dotted lines in (b). Stable (unstable) solutions are shown with solid (dotted) lines. The bifurcation point is $(u_{0}^{*},0)$. (b): Coefficient $p$ ($q$) as a function of $K_{u}$ shown as a solid (dashed) black line.
Proposition 1 uncovers the key role of $K_{u}$ in tuning opinion formation: $K_{u}$ controls the supercritical versus subcritical nature of opinion formation and increasing $K_{u}$ increases opinion strength at non-neutral bifurcation branches (Fig. 2). The bifurcation diagrams in Fig. 2a are plots of the equilibrium solutions of (2), i.e., the solutions of $\dot{z}\!=\!0$, as a function of bifurcation parameter $u_{0}$, for five different values of $K_{u}$ distinguished by color. The singularity at $(u_{0},z)\!=\!(u_{0}^{*},0)$ is a pitchfork bifurcation: blue, gray, and green lines show supercritical, quintic, and subcritical solutions, respectively. The corresponding values of $K_{u}$ are indicated in Fig. 2b. As $K_{u}$ increases from 00, the supercritical pitchfork grows wider (increasing opinion strength) as $p$ grows less negative, becomes quintic when $p\!=\!0$, and becomes a subcritical pitchfork when $p\!>\!0$. For the subcritical pitchforks in Fig. 2a, two stable non-neutral solutions appear for $u_{0}\!<\!u_{0}^{*}$, through so-called saddle-node bifurcations. As $p$ grows more positive, these solutions emerge for smaller values of $u_{0}$ and increase in magnitude, reflecting increasing opinion strength.

### v3

# Spiking Nonlinear Opinion Dynamics (S-NOD) for Agile Decision-Making

@@ Abstract
We present, analyze, and illustrate a first-of-its-kind model of two-dimensional excitable (spiking) dynamics for decision-making over two options. The model, Spiking Nonlinear Opinion Dynamics (S-NOD), provides superior agility, characterized by fast, flexible, and adaptive response to rapid and unpredictable changes in context, environment, or information received about available options. S-NOD derives through the introduction of a single extra term to the previously presented Nonlinear Opinion Dynamics (NOD) for fast and flexible multi-agent decision-making behavior. The extra term is inspired by the fast-positive, slow-negative mixed-feedback structure of excitable systems. The agile behaviors brought about by the new excitable nature of decision-making driven by S-NOD are analyzed in a general setting and illustrated in an application to multi-robot navigation around human movers.
@@ I Introduction
The fast, flexible, and adaptive behavior observed in biology owes much to the excitable (spiking) nature of cellular signaling [1, 2, 3, 4]. Models of excitability represent the analog molecular and/or biophysical processes that produce spikes in response to stimuli. These models inherit the adaptive behavior of analog systems and the reliability of digital systems, foundational to spiking control systems [5] and neuromorphic engineering [6]. However, existing models describe single-input/single-output [7] spike-based signal processing. This is spiking activity that can only encode single-option decisions: to spike or not spike, as determined by the input signal pushing the system toward its excitability threshold [8]. This limits the use of these models in studying and designing spiking decision-making over multiple options as observed, e.g., in cortical column visual orientation selectivity [9] or in sensorimotor decision-making [10].
We present a generalized model of excitable (spiking) dynamics that allows for fast, flexible, and adaptive decision-making over multiple options. In this paper we focus on two-option spiking in a two-dimensional, two-timescale model and we use “agile” to mean “fast, flexible, and adaptive.” To the best of our knowledge, this is the first such model to generalize spiking to more than one option, i.e., spiking that can occur in any of the multiple directions corresponding to the multiple options available to the excitable decision-maker. We call our model Spiking Nonlinear Opinion Dynamics (S-NOD) as it derives from the introduction of an extra term in the Nonlinear Opinion Dynamics (NOD) model of [11, 12] that makes NOD excitable, i.e., spiking.
@@
NOD model the time evolution of opinions of a group of agents engaged in a collective decision-making process over a set of options. The derivation of NOD was tailored to model and study the principles of fast and flexible decision-making in biological collectives [13, 14] and to use these principles to design fast and flexible decision-making in built collectives [11, 15, 16]. NOD exhibits a mixed-feedback [17, 8] structure: opinion formation arises from the balance of a negative feedback loop that regulates agent opinions to a neutral solution and positive feedback loops (at the single-agent and network levels) that destabilize the neutral solution and trigger nonlinear opinion formation. Decision-making driven by NOD [12] is fast because it can diverge quickly from indecision even in the absence of informative inputs about the options. It is flexible because the sensitivity of opinion formation to informative inputs is tunable. Both speed and flexibility are determined by a tunable threshold for opinion formation where negative and positive feedback are perfectly balanced and the dynamics become singular.
@@
Figure 1: (a): Trajectory of a robot controlled with NOD (S-NOD) is shown with a blue (pink) line as it navigates towards a goal (star) in the presence of an oncoming human mover (black). The NOD robot experiences a collision; the S-NOD robot does not. (b): Opinion $z$ of the robot over time $t$. Circles denote matching time points along trajectories and opinions.
Figures 1, 4, 5, and 6 animated at https://spikingNOD.github.io.
@@
To derive S-NOD, we introduce in NOD a slow regulation term inspired by the dynamics of excitable (spiking) signal processing systems. The resulting excitable dynamics give S-NOD its superior agility in decision-making. Where NOD allows for a fast decision, S-NOD allows for autonomous fast sequential decision-making, not requiring any ad-hoc reset of the model state once a decision is made. Where NOD is flexible, S-NOD is flexible and capable of fast “changes of mind” and adaptive responses when the information about the options changes rapidly and unexpectedly. Further, S-NOD provides on-demand (event-based) opinion formation in the sense that large opinions are formed sparsely in time in the form of “decision events” and only when context requires it. This makes S-NOD efficient. The agility of S-NOD is illustrated in Fig. 1 in the context of a robot navigating around a human mover as studied in [18].
@@
Our major contributions in this paper are the presentation, analysis, and illustration of a first-of-its-kind model of two-dimensional excitable (spiking) dynamics for decision-making over two options, which provides superior agility (fast, flexible and, adaptive behavior), especially important in changing contexts. Also, in Section II, we present a new analysis of the singularity in NOD for a single decision-making agent and two options. We prove how a single feedback gain $K_{u}$ tunes opinion formation. We show that when $K_{u}$ gets too large, an opinion can become so robust that it will not change quickly enough if a new input arrives in favor of the alternative option. In Section III, we present S-NOD for a single agent and two options. We show the existence of the spiking limit cycles associated with excitable behavior. We show geometrically the agility in behavior. We generalize S-NOD to a network of multiple agents and apply it to a social robot navigation problem in Section IV.
II Fast and Flexible Decision-Making: NOD
@@
We recall NOD [11, 12] in Section II-A for a single agent evolving continuously over time its real-valued opinion about two mutually exclusive options with possible input present. In Section II-B we analyze stability of the neutral opinion solution and prove conditions on feedback gain $K_{u}$ that determine the type of singularity (type of pitchfork bifurcation) in the dynamics. We show that, by shaping bifurcation branches, $K_{u}$ tunes opinion formation. In Section II-C, we show limits on tunability of NOD that sacrifice agility in decision-making, motivating S-NOD in Section III.
II-A NOD for Single Decision-Maker and Two Options
@@
We let an agent represent a single decision-maker. Let $z(t)\in\mathbb{R}$ define the agent’s opinion at time $t$ about two mutually exclusive options. The more positive (negative) is $z$, the more the agent favors (disfavors) option 1 and disfavors (favors) option 2. When $z\!=\!0$, the agent is neutral about the two options, i.e., in a state of indecision. Let $u(t)\!\geq\!0$ define the attention of the agent at time $t$ to its observations; $u$ is implemented as a gain in the dynamics. Let $b(t)\in\mathbb{R}$ define an input signal at time $t$ that represents external stimulus and/or internal bias. When $b(t)\!>\!0$ ($b(t)\!<\!0$), it provides information (evidence) in favor of option 1 (option 2).
@@
Decision-making variables $z$, $u$ evolve in continuous time $t$ according to the following NOD, adapted from [11, 12]:
$\tau_{z}\dot{z}$ $=-d\,z+\tanh\big{(}u\left(az\right)+b\big{)}$ (1a)
@@
where ${\dot{z}}:=\!dz/dt$. $\tau_{z}\!>\!0$ is a time constant, and damping coefficient $d\!>\!0$ weights the negative feedback on $z$ that regulates to the neutral solution $z\!=\!0$. The second term in (1a) provides a nonlinear positive feedback on $z$ with weight given by the product of $u$ and amplification coefficient $a\!>\!0$, plus the effects of $b$. The saturation nonlinearity given by the $\tanh$ function enables fast-and-flexible decision-making through opinion-forming bifurcations [11, 12]. The positive feedback gain is state-dependent according to (1b) and grows with $z^{2}$. Hence, small deviations from the neutral solution ($z\!=\!0$) in response to small inputs leave attention low and do not trigger large, nonlinear opinion formation. Large enough deviations from the neutral solution in response to large enough inputs cause a sharp increase in attention and trigger large, nonlinear opinion formation. The resulting implicit threshold distinguishing small and large inputs is tuned by $u_{0}$ and $K_{u}$.
II-B Analysis of Single-Agent, Two-Option NOD
We study the dynamics and stability of solutions of system (1a)-(1b) using bifurcation theory. A local bifurcation refers to a change in number and/or stability of equilibrium solutions of a nonlinear dynamical system as a (bifurcation) parameter is changed. The state and parameter values at which this change occurs is the bifurcation point. At a bifurcation point, one or more eigenvalues of the Jacobian of the model must have zero real part [19, 20], i.e., a bifurcation point is a singularity of the model vector field.
Our main interest is in the pitchfork bifurcations. There are two generic types of pitchforks. A supercritical pitchfork bifurcation describes how one stable solution becomes unstable and two stable solutions emerge as the bifurcation parameter increases. A subcritical pitchfork bifurcation describes how two unstable solutions disappear and one stable solution becomes unstable as the bifurcation parameter increases.
@@ II-B Analysis of Single-Agent, Two-Option NOD
$\tau_{z}\dot{z}=-d\,z+\tanh\Big{(}(u_{0}+K_{u}z^{2})\cdot\left(az\right)+b\Big{)}.$ (2)
@@
$\dot{z}=\frac{1}{\tau_{z}}\Bigg{(}\left(a\,u_{0}-d\right)z+a\left(K_{u}-\frac{a^{2}u_{0}^{3}}{3}\right)z^{3}\\ +a^{3}u_{0}^{2}\left(\frac{2a^{2}u_{0}^{3}}{15}-K_{u}\right)z^{5}\Bigg{)}+\mathcal{O}(z^{7}).$ (3)
A singularity exists at $(u_{0},z)\!=\!(u_{0}^{*},0)$, with $u_{0}^{*}\!=\!\frac{d}{a}$. The solution $z\!=\!0$ is stable (unstable) when $u_{0}\!<\!u_{0}^{*}$ ($u_{0}\!>\!u_{0}^{*}$).
@@
When $z\!=\!0$ and $b\!=\!0$, the right hand side of (2) is zero thus $z\!=\!0$ is always an equilibrium. We expand (2) with $b\!=\!0$ about $z\!=\!0$. The Taylor expansion of the hyperbolic tangent is $\tanh(w)\!=\!w-w^{3}/3+2w^{5}/15+\mathcal{O}(w^{7})$. Substituting this into (2) yields (3). The Jacobian $J(z)\!=\!\frac{d\dot{z}}{dz}$ of (3) evaluated at $z\!=\!0$ is $J(0)\!=\!(a\,u_{0}-d)/\tau_{z}$, which is singular when $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{a}$. When $u\!<\!u_{0}^{*}$, $J(0)\!<\!0$ thus $z\!=\!0$ is exponentially stable. When $u\!>\!u_{0}^{*}$, $J(0)\!>\!0$ thus $z\!=\!0$ is unstable. ∎
@@
($K_{u}$ determines singularity type): Let $b\!=\!0$, $u_{0}^{*}\!=\!\frac{d}{a}$. The singularity of dynamics (2) at $(u,z)\!=\!(u_{0}^{*},0)$ as proved in Lemma 1 corresponds to a supercritical pitchfork bifurcation for $K_{u}\!<\!\frac{d^{3}}{3a}$, a quintic pitchfork bifurcation for $K_{u}\!=\!\frac{d^{3}}{3a}$, and a subcritical pitchfork bifurcation for $K_{u}\!>\!\frac{d^{3}}{3a}$.
@@
Denote $p\!=\!(K_{u}\!-\!\frac{d^{3}}{3a})$ and $q\!=\!(\frac{2d^{3}}{15a}\!-\!K_{u})$ as the coefficients of $az^{3}/\tau_{z}$ and $(a^{3}u_{0}^{2})z^{5}/\tau_{z}$ in (3) resp. at $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{a}$. When $K_{u}\!<\!\frac{d^{3}}{3a}$ ($K_{u}\!>\!\frac{d^{3}}{3a}$), then $p\!<\!0$ ($p\!>\!0$) and (3) is the normal form of the supercritical (subcritical) pitchfork bifurcation [19]. When $K_{u}\!=\!\frac{d^{3}}{3a}$, then $p\!=\!0$, $q\!<\!0$ and (3) is the normal form of the quintic pitchfork, by recognition problem [19, Prop. VI.2.14] and its $Z_{2}$-symmetric universal unfolding [19, Prop. VI.3.4; Fig. VI.3.3]. ∎
Figure 2: The effect of $K_{u}$ on the bifurcation diagram of (2) and the cubic and quintic terms of (3). (a): Bifurcation diagrams of NOD (2) with $K_{u}$ values corresponding to the vertical dashed lines in (b). Stable (unstable) solutions are shown with solid (dotted) lines. The bifurcation point is $(u_{0}^{*},0)$. (b): Coefficient $p$ ($q$) as a function of $K_{u}$ shown as a solid (dashed) black line.
Proposition 1 uncovers the key role of $K_{u}$ in tuning opinion formation (Fig. 2): (i) $K_{u}$ controls the supercritical vs. subcritical nature of opinion formation, and (ii) increasing $K_{u}$ increases opinion strength of non-neutral solutions. Bifurcation diagrams in Fig. 2a plot equilibrium solutions of (2), i.e., solutions of $\dot{z}\!=\!0$, as a function of bifurcation parameter $u_{0}$ for different values of $K_{u}$ (Fig. 2b). The singularity at $(u_{0},z)\!=\!(u_{0}^{*},0)$ is a pitchfork bifurcation: blue, gray, green lines show supercritical, quintic, subcritical solutions, respectively. For all $K_{u}$: when $u_{0}\!<\!u_{0}^{*}$, the neutral solution is stable; when $u_{0}\!\!>\!\!u_{0}^{*}$, the neutral solution is unstable and there is a bistable symmetric pair of solutions. When $K_{u}\!\!<\!\!\frac{d^{3}}{3a}$, the pitchfork is supercritical: there are no other solutions. When $K_{u}\!\!>\!\!\frac{d^{3}}{3a}$, the pitchfork is subcritical: two stable non-neutral solutions appear for $u_{0}\!<\!u_{0}^{*}$, through saddle-node bifurcations. As $K_{u}$ grows more positive, these solutions emerge for smaller values of $u_{0}$ and increase in magnitude, reflecting increasing opinion strength.
Figure 3: Opinion solutions of NOD and S-NOD over time and associated bifurcation diagrams. (A): Trajectories of NOD (2) and S-NOD (4) for initial condition $(z,u_{0})|_{t=0}=(0.01,0.9)$ for two $K_{u}$ values. Input signal $b$ is also shown over time. (B): Bifurcation diagrams of (2) for the two values of $K_{u}$, showing the solutions in A. moving from initial state to steady state.